OpenStudy (anonymous):

Is the answer to this x=-21 3-6/7x=21

OpenStudy (anonymous):

(3-6)/(7x) = 21 if you place -21 you will get a really large negative number in the denominator compared to the numerator which will result in a extremely small number

OpenStudy (anonymous):

It is 3- 6 over 7x

OpenStudy (anonymous):

3-(6/7x) will also do the same but you are subtracting 3 by a small number

OpenStudy (anonymous):

o

OpenStudy (anonymous):

if you multiply both sides by x you can see that you will get (3-6)/7 = 21x

OpenStudy (anonymous):

you can simplify 3-6 to -3 so you will result in -3/7 = 21x

OpenStudy (anonymous):

then you can divide both sides by 21 to get -3/(7*21) = x and that is your answer

OpenStudy (anonymous):

to check you can place your answer of x into the original equation and it should equal 21, the right hand side of the equation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

should it be a negative number? -21?